The doctor order Atropine gr. 1/200 amd available is atropine gr. 1/150 in eml how many ml will you give. Please show response
Answer #1

THE  DOCTOR ORDER IS -1/200

THE AVAILABLE AMOUNT -1/150 IN ML

SO HOWMANY ML WILL GIVE

doctor amount / theavailable amount

1/200 x 150=150/200=.75 ML